OpenStudy (unklerhaukus):

Show that: \[\newcommand \dd {\,\mathrm d} \boxed{ \int x^2e^{ax}\dd x = \left(\frac{x^2}a - \frac{2x}{a^2} + \frac2{a^3}\right)e^{ax} }\]

myininaya (myininaya):

what about integration by parts

OpenStudy (unklerhaukus):

\[ \int x^2e^{ax}\dd x = \frac{x^2}ae^{ax}-\frac{2}a\int xe^{ax}\dd x \]

OpenStudy (unklerhaukus):

\[=\frac{x^2}ae^{ax}-\frac2a\left[\frac xae^{ax}-\frac1a\int e^{ax}\dd x\right]\]

OpenStudy (unklerhaukus):

\[=\frac{x^2}ae^{ax}-\frac{2x}{a^2}e^{ax}+\frac2{a^2}\int e^{ax}\dd x\\ =\frac{x^2}ae^{ax}-\frac{2x}{a^2}e^{ax}+\frac2{a^3}e^{ax}+C\\ =\left(\frac{x^2}{a}-\frac{2x}{a^2}+\frac2{a^3}\right)e^{ax}+C\]

myininaya (myininaya):

you can also differentiate both sides to prove that one thing is equal to the other

ganeshie8 (ganeshie8):

we may try differentiation under the integral : \[\int x^2e^{ax}\,dx=\int \dfrac{\partial^2}{\partial a^2}e^{ax}\,dx\]

myininaya (myininaya):

\[\text{we can say } \int\limits f(x) dx=g(x) \text{ by verifying } f(x)=g'(x)\]

ganeshie8 (ganeshie8):

More generally : \[\int x^ne^{ax}\,dx=\int \dfrac{\partial^n}{\partial a^n}e^{ax}\,dx = \dfrac{\partial^n}{\partial a^n} \int e^{ax}\, dx = \dfrac{\partial^n}{\partial a^n} \frac{e^{ax}}{a}\]

OpenStudy (unklerhaukus):

\[\frac{\dd}{\dd x}\left(\frac{x^2}a - \frac{2x}{a^2} + \frac2{a^3}\right)e^{ax}\\ \qquad=\left[\left(\frac{2x}a - \frac{2}{a^2}\right)+\left({x^2} - \frac{2x}{a} + \frac2{a^2}\right)\right]e^{ax}\\ \qquad =x^2e^{ax}\\[4ex] \qquad \Downarrow\\[2ex] \int x^2e^{ax}\dd x = \left(\frac{x^2}a - \frac{2x}{a^2} + \frac2{a^3}\right)e^{ax} \]

OpenStudy (unklerhaukus):

\[\begin{align} \int x^2e^{ax}\dd x &=\int \frac{\partial^2}{\partial a^2}e^{ax}\dd x\\ &= \dfrac{\dd ^2}{\dd a^2} \int e^{ax}\dd x\\ &= \dfrac{\dd^2}{\dd a^2}\left(\frac{e^{ax}}a\right)\\ &=\frac{\dd }{\dd a}\left(\frac{ax-1}{a^2}e^{ax}\right)\\ &=\frac{a^2\big(x+x(ax-1)\big)-2a\big(ax-1\big)}{a^4}e^{ax}\\ &=\frac{\Big(a^2x+a^3x^2-a^2x\Big)-\Big(2a^2x-2a\Big)}{a^4}e^{ax}\\ &=\frac{a^3x^2-2a^2x+2a}{a^4}e^{ax}\\ &=\left(\frac{x^2}a-\frac{2x}{a^2}+\frac{2}{a^3}\right)e^{ax}\\ \end{align}\]

OpenStudy (unklerhaukus):

Thank you @myininaya Thank you @ganeshie8